# Positive Semidefinite Supermartingales and Randomized Matrix Concentration Inequalities

### v3

@@ 6.1 Comparison with a Single Observation
When only $n=1$ observation is revealed, $L_{1}^{\textsf{SN}}=\mathrm{tr}(\mathbf{Y}_{1}^{\textsf{SN}})$. We shall see that the comparison depends heavily on the rejection threshold matrix $\mathbf{A}$ that Matrix picks beforehand.
@@
with $\varepsilon\in(0,\alpha-\lambda_{d}^{-1})$. The condition $\lambda_{d}>1/\alpha$ here is strictly weaker than the condition under which Scalar rejects: $\lambda_{1}+\dots+\lambda_{d}=d/\alpha$.
Indeed, as we can see from these examples with $n=1$ (thus we have only compared the matrix Markov inequality with the scalar Markov inequality on the trace), for Matrix to outpower Scalar, apart from prior knowledge, it is favorable that $\lambda_{\max}(\mathbf{Y}_{1}^{\mathsf{SN}})$ is an outlier in the spectrum of $\mathbf{Y}_{1}^{\mathsf{SN}}$, i.e., ${\lambda_{\max}(\mathbf{Y}_{1}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{1}^{\mathsf{SN}})}$ is close to 1. This is exactly what we would predict from the discussion on (UMMI) early on, where we mention that (UMMI) is tighter when the random matrix has (approximately) a lower rank. In the very next subsection, we shall see that whereas the condition ${\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}\approx 1$ seems artificial and contrived here with $n=1$, it becomes an almost certain phenomenon with large $n$.
@@ 6.2 Comparison with a Sequence of Observations
The scalar component process $Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})$ grows exponentially fast if $\pi_{k}$ has mean $\lambda_{k}\neq 0$, and shrinks to 0 if $\lambda_{k}=0$; moreover, the rate in which $Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})$ grows depends exponentially on the size $|\lambda_{k}|$. Consequently, as long as there is any mean disparity between the component distributions $\pi_{1},\dots,\pi_{d}$, say $|\lambda_{1}|,\dots,|\lambda_{d-1}|<|\lambda_{d}|$, in the long run the process $\mathbf{Y}_{n}^{\textsf{SN}}$ will be approximately low-rank with ${\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}\to 1$, as the component $Y_{n}^{\textsf{SN}}(\lambda_{d1},\dots,\lambda_{dn})\mathbf{P}_{d}$ will exponentially outgrow the rest. In this case if Matrix correctly anticipates the largest component $\mathbf{P}_{d}$, by picking again the threshold matrix
@@
We experiment with the following setup. We pick an orthonormal basis in $d=3$ dimension: $\mathbf{u}_{1}=(1/\sqrt{2},1/\sqrt{2},0)^{\mathsf{T}}$, $\mathbf{u}_{2}=(0,0,1)^{\mathsf{T}}$, $\mathbf{u}_{3}=(1/\sqrt{2},-1/\sqrt{2},0)^{\mathsf{T}}$, and let $\mathbf{P}_{1}=\mathbf{u}_{1}\mathbf{u}_{1}^{\mathsf{T}}$, $\mathbf{P}_{2}=\mathbf{u}_{2}\mathbf{u}_{2}^{\mathsf{T}}$, $\mathbf{P}_{3}=\mathbf{u}_{3}\mathbf{u}_{3}^{\mathsf{T}}$. The spectral distributions are set as $\pi_{1}=\mathcal{N}(0.1,1)$, $\pi_{2}=\mathcal{N}(-0.2,1)$, and $\pi_{3}=\mathcal{N}(0.3,1)$. We consider the following three sequential tests: smart Matrix, one that follows the recipe (99) with $\varepsilon=\alpha/20$; naive Matrix, one that is agnostic about the real distribution and picks $\mathbf{A}=(d/\alpha)\mathbf{I}$; and Scalar, all with the parameters $\gamma_{n}=n^{-1/2}$ and $\alpha=0.05$. We record the times at which these tests correctly reject the null, as well as the evolution of ${\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}$. The findings are displayed in Figure 1: smart Matrix indeed outpowers the other two, whereas naive Matrix and Scalar are similar in power, explained by the quick convergence of ${\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}$ to 1.
Figure 1: Empirical findings with commuting i.i.d. random matrices. Left. Comparison of the cumulative distribution functions of rejection times of the sequential tests corresponding to the matrix- and scalar-valued test processes under the same alternative distribution. Right. Evolution of the spectral distribution of the matrix-valued test process under the alternative.
@@
which indeed satisfies $(\mathrm{tr}\mathbf{\Sigma})^{2}+2\mathrm{tr}(\mathbf{\Sigma}^{2})=\frac{61}{144}<1$. The three sequential tests as before are compared in Figure 2: the smart Matrix that follows (101) with $\varepsilon=\alpha/20$; the naive Matrix with $\mathbf{A}=(d/\alpha)\mathbf{I}$; Scalar, all with $\gamma_{n}=n^{-1/2}$ and $\alpha=0.05$. Besides the unsurprising fact that the smart Matrix is again the most powerful, rejecting the null earlier than the other two, we also observe the phenomenon that, while in the $n=1$ case, the matrix-valued test process with “naive” $\mathbf{A}=(d/\alpha)\mathbf{I}$ is always less powerful than the scalar-valued test process, this is reversed under non-commutativity with large $n$ as naive Matrix rejects slightly earlier than Scalar in our simulation. As also shown in Figure 2, the combination of two auxiliary findings contributes to this phenomenon: ${\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}\to 1$ even when observations do not commute; and, while $L_{n}^{\textsf{SN}}=\mathrm{tr}(\mathbf{Y}_{n}^{\textsf{SN}})$ in all our previous examples, now $\mathrm{tr}(\mathbf{Y}_{n}^{\textsf{SN}})$ grows slightly faster than $L_{n}^{\textsf{SN}}$. It takes further theoretical investigations to understand why this happens (either in this particular example or in general), but to some extent it hints nonetheless at the possibility that in practice where the effect size is small (requiring large $n$) and matrix observations are non-commuting, there is an inherent advantage using the matrix-valued test process we propose.
Figure 2: Empirical findings with non-commuting i.i.d. random matrices for covariance testing. Left. Comparison of the cumulative distribution functions of rejection times of the sequential tests corresponding to the matrix- and scalar-valued test processes under the same alternative distribution. Middle. Evolution of the spectral distribution of the matrix-valued test process under the alternative. Right. Comparison between the trace of the matrix-valued test process, and the scalar-valued test process under the same alternative distribution.
@@ 7 Conclusion
We prove in this paper several novel concentration inequalities for random symmetric matrices. Some of these bounds are tighter than previous ones in the literature via randomization or Loewner order, others enjoy the virtues of validity at stopping times and/or the more general applicability to non-i.i.d. exchangeable matrices. These results shall find applications in, for example, covariance estimation and testing.
We also leave open some questions for future studies to address. These include a deeper understanding of the matrix randomizer $\mathbf{U}$ in our paper, which belongs to the “trace super-uniform” distribution family; tighter $p$-Chebyshev inequalities; further investigation into matrix-valued e-processes; as well as the theoretical explanation of our comparison between matrix-valued test supermartingales and scalar-valued test supermartingales when testing sequentially the same null hypotheses. Our experiments show that matrix-valued test processes are often more powerful in these sequential tests.
@@ Acknowledgement
The authors thank Joel A. Tropp, Lester Mackey, and Tudor Manole for suggestions and discussions, and acknowledge support from NSF grant DMS-2310718.
Appendix A $p$-Chebyshev Inequalities
A.1 The Scalar (and Vector) $p$-Chebyshev Inequality
The following statement shows that for i.i.d. random scalars, the central $p$th moment of the sum grows linearly in the number of scalars just like the variance.
Proposition A.1 (Contraction of the scalar $p$th moment, Lemma 4.2 of Cherapanamjeri et al. [2022]).
Suppose $1\leqslant p\leqslant 2$. Let $X_{1},\dots,X_{n}$ be i.i.d. scalar with mean $\mu=\mathbb{E}X_{1}$ and central $p$th moment $v_{p}=\mathbb{E}|X_{1}-\mu|^{p}$. Then, $\mathbb{E}|X_{1}+\dots+X_{n}-n\mu|^{p}\leqslant 2^{2-p}nv_{p}$.
Wang et al. [2021] generalize the statement to vectors under the martingale difference assumption.
Proposition A.2 (Concentration of the vector $p$th moment, Lemma 7 of Wang et al. [2021]).
Suppose $1\leqslant p\leqslant 2$. Let $\mathbf{x}_{1},\dots,\mathbf{x}_{n}$ be $d$-dimensional random vectors adapted to a filtration $\mathcal{F}_{0},\mathcal{F}_{1},\dots,\mathcal{F}_{n}$ with conditional mean $\mu=\mathbb{E}(\mathbf{x}_{i}|\mathcal{F}_{i-1})$ and conditional central $p$th moment upper bound $v_{p}\geqslant\mathbb{E}(\|\mathbf{x}_{i}-\mu\|^{p}|\mathcal{F}_{i-1})$. Then, $\mathbb{E}\|\mathbf{x}_{1}+\dots+\mathbf{x}_{n}-n\mu\|^{p}\leqslant 2^{2-p}d^{1-p/2}nv_{p}$.
@@ A.1 The Scalar (and Vector) $p$-Chebyshev Inequality
Corollary A.3 (Vector $p$-Chebyshev Inequality).
Suppose $\mathbf{x}_{1},\dots,\mathbf{x}_{n}$ satisfy the same assumptions as in Proposition A.2. Then, for any $a>0$,
$\mathbb{P}\left(\left\|\overline{\mathbf{x}}_{n}-\mu\right\|\geqslant a\right)\leqslant\frac{2^{2-p}d^{1-p}v_{p}}{n^{p-1}a^{p}}.$ (103)
A.2 Matrix $p$-Chebyshev Inequalities
The following conjecture is an ideal extension of Propositions A.2 and A.3.
Conjecture A.4 (Matrix $p$-Chebyshev Inequality).
Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be i.i.d. random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{1}=\mathbf{M}$ and $p$th central moment matrix $\mathbf{V}_{p}:=\mathbb{E}(\operatorname{abs}(\mathbf{X}_{1}-\mathbf{M}))^{p}$. Then, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$, there exists a function $f:[1,2]\times\mathbb{N}\to(0,\infty)$ that grows sublinearly in its second argument, and bounded in its first argument, such that the $p$th central moment contracts,
$\mathbb{E}(\operatorname{abs}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n}-n\mathbf{M})^{p})\preceq n\cdot f(p,d)\cdot\mathbf{V}_{p};$ (104)
and consequently, for $\overline{\mathbf{X}}_{n}=\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$,
$\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant n^{1-p}\cdot f(p,d)\cdot\mathrm{tr}(\mathbf{V}_{p}\mathbf{A}^{-p}).$ (105)
If it holds, the $\mathbf{A}$ above can be randomized into $(\mathbf{A}^{p/2}\mathbf{U}\mathbf{A}^{p/2})^{1/p}$ as well.
We do have the following heavily scalarized and dimension-dependent bound that scales correctly w.r.t. the sample size $n$. We find this bound trivial as it regards the matrix as a mere collection of $d^{2}$ scalars and is strictly looser than the “standard” vector bound (Proposition A.2) applied to the flattened matrix.
Proposition A.5 (Matrix $p$-Chebyshev Spectral Inequality).
Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be i.i.d. random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{1}=\mathbf{M}$ and $p$th spectral central moment $v_{p}:=\mathbb{E}\|\mathbf{X}_{1}-\mathbf{M}\|^{p}$. Then,
$\mathbb{E}\|\mathbf{X}_{1}+\dots+\mathbf{X}_{n}-n\mathbf{M}\|^{p}\leqslant 2^{2-p}d^{2-p/2}\cdot n\cdot v_{p}.$ (106)
Proof.
Without loss of generality suppose $\mathbf{M}=0$. Denote by $\|\cdot\|_{L_{p}}$ the entrywise matrix $p$-norm and $\|\cdot\|_{\mathrm{F}}=\|\cdot\|_{L_{2}}$ the Frobenius norm. Recall that $\|\mathbf{A}\|\leqslant\|\mathbf{A}\|_{\mathrm{F}}\leqslant\sqrt{d}\|\mathbf{A}\|$ and $\|\mathbf{A}\|_{\mathrm{F}}\leqslant\|\mathbf{A}\|_{L_{p}}\leqslant d^{\frac{2}{p}-1}\|\mathbf{A}\|_{\mathrm{F}}$. Then, $\mathbb{E}\|\mathbf{X}_{1}+\dots+\mathbf{X}_{n}\|^{p}$ is upper bounded by
$\mathbb{E}\|\mathbf{X}_{1}+\dots+\mathbf{X}_{n}\|_{L_{p}}^{p}=\mathbb{E}\left[\sum_{1\leqslant j,k\leqslant d}\left|\sum_{i=1}^{n}X_{i(jk)}\right|^{p}\right]=\sum_{1\leqslant j,k\leqslant d}\mathbb{E}\left|\sum_{i=1}^{n}X_{i(jk)}\right|^{p}$ (107)
$\stackrel{{\scriptstyle(*)}}{{\leqslant}}\,$ $\sum_{1\leqslant j,k\leqslant d}2^{2-p}n\mathbb{E}\left|X_{1(jk)}\right|^{p}=2^{2-p}n\mathbb{E}\|\mathbf{X}_{1}\|_{L_{p}}^{p}\leqslant 2^{2-p}d^{2-p}\cdot n\cdot\mathbb{E}\|\mathbf{X}_{1}\|^{p}_{\mathrm{F}}$ (108)
[…]
where the inequality (*)(*)( * ) comes from the contraction of scalar $p$th moment (Proposition A.1). ∎
Appendix B Additional Lemmas and Omitted Proofs
B.1 Proofs of Statements in Table 1
Proof.
For Rademacher, the bound $\mathbf{G}(\gamma)\preceq\exp(\gamma^{2}\mathbf{C}^{2}/2)$ is proved by Tropp [2015, Lemma 4.6.3] via a direct Taylor expansion argument. Since the matrix log is monotone (Fact 2), $\log\mathbf{G}(\gamma/n)\preceq\gamma^{2}\mathbf{C}^{2}/2n^{2}$. Finally, invoking the monotonicity of $\mathbf{A}\mapsto\mathrm{tr}\exp\mathbf{A}$ (Fact 3) we see that $\mathrm{tr}(\mathrm{e}^{n\log\mathbf{G}(\gamma/n)})\leqslant\mathrm{tr}(\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2n})$. Bennett I follows from a similar reasoning applied to the Bennett MGF bound [Tropp, 2012, Lemma 6.7]. For Bennett II, we can apply the variance bound at last: $\mathrm{tr}(\mathrm{e}^{n(\mathrm{e}^{\gamma/n}-\gamma/n-1)\mathbf{Var}(\mathbf{X}_{1})})\leqslant\mathrm{tr}(\mathrm{e}^{n(\mathrm{e}^{\gamma/n}-\gamma/n-1)\mathbf{V}})$. The uni-Gaussian MGF bound is also proved by Tropp [2015, Lemma 4.6.2] via a direct Taylor expansion argument. For symmetric Hoeffding, introducing an independent $R\sim\operatorname{rad}_{1/2}$,
$\log\mathbf{G}(\gamma)=\log\mathbb{E}\mathrm{e}^{\gamma(\mathbf{X}_{1}-\mathbf{M})}=\log\mathbb{E}\mathrm{e}^{\gamma R(\mathbf{X}_{1}-\mathbf{M})}\preceq\gamma^{2}\mathbf{B}/2$ (110)
due to [Tropp, 2012, Lemma 7.7], leading to the corresponding upper bound on $\mathrm{tr}(\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2n})$ as before. ∎
B.2 Dominated Integrability Lemma

### v4

@@ 6.1 Comparison with a Single Observation
When only $n=1$ observation is revealed, $L_{1}^{\textsf{SN}}=\mathrm{tr}(\mathbf{Y}_{1}^{\textsf{SN}})$. We shall see that the comparison depends heavily on the rejection threshold matrix $\mathbf{A}$ that Matrix picks beforehand, as rejection happens on the complements of the “concentration regions” discussed in Section 5.1. The following are direct corollaries of Proposition 5.1.
@@
with $\varepsilon\in(0,\alpha-\lambda_{d}^{-1})$. The condition $\lambda_{d}>1/\alpha$ here is strictly weaker than the condition under which Scalar rejects: $\lambda_{1}+\dots+\lambda_{d}\geqslant d/\alpha$.
Indeed, as we can see from these examples with $n=1$ (thus we have only compared the matrix Markov inequality with the scalar Markov inequality on the trace), for Matrix to outpower Scalar, apart from prior knowledge, it is favorable that $\lambda_{\max}(\mathbf{Y}_{1}^{\mathsf{SN}})$ is an outlier in the spectrum of $\mathbf{Y}_{1}^{\mathsf{SN}}$, i.e., its effective rank $\operatorname{r}(\mathbf{Y}_{1}^{\mathsf{SN}})={\mathrm{tr}(\mathbf{Y}_{1}^{\mathsf{SN}})}/{\lambda_{\max}(\mathbf{Y}_{1}^{\mathsf{SN}})}$ is close to 1. This is exactly what we would predict from the discussion in Section 5.1, where we mention that (MMI) is tighter when the rejection threshold matrix $\mathbf{A}$ (thus also the random matrix itself) has a lower effective rank. In the very next subsection, we shall see that whereas the low-rank condition $\operatorname{r}(\mathbf{Y}_{n}^{\mathsf{SN}})\approx 1$ seems artificial and contrived here with $n=1$, it becomes an almost certain phenomenon with large $n$.
@@ 6.2 Comparison with a Sequence of Observations
The scalar component process $Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})$ grows exponentially fast if $\pi_{k}$ has mean $\lambda_{k}\neq 0$, and shrinks to 0 if $\lambda_{k}=0$; moreover, the rate in which $Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})$ grows depends exponentially on the size $|\lambda_{k}|$. Consequently, as long as there is any mean disparity between the component distributions $\pi_{1},\dots,\pi_{d}$, say $|\lambda_{1}|,\dots,|\lambda_{d-1}|<|\lambda_{d}|$, in the long run the process $\mathbf{Y}_{n}^{\textsf{SN}}$ will have a low effective rank with $\operatorname{r}(\mathbf{Y}_{n}^{\mathsf{SN}})={\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}\to 1$, as the component $Y_{n}^{\textsf{SN}}(\lambda_{d1},\dots,\lambda_{dn})\mathbf{P}_{d}$ will exponentially outgrow the rest. In this case if Matrix correctly anticipates the largest component $\mathbf{P}_{d}$, by picking again the threshold matrix
@@
We experiment with the following setup. We pick an orthonormal basis in $d=3$ dimension: $\mathbf{u}_{1}=(1/\sqrt{2},1/\sqrt{2},0)^{\mathsf{T}}$, $\mathbf{u}_{2}=(0,0,1)^{\mathsf{T}}$, $\mathbf{u}_{3}=(1/\sqrt{2},-1/\sqrt{2},0)^{\mathsf{T}}$, and let $\mathbf{P}_{1}=\mathbf{u}_{1}\mathbf{u}_{1}^{\mathsf{T}}$, $\mathbf{P}_{2}=\mathbf{u}_{2}\mathbf{u}_{2}^{\mathsf{T}}$, $\mathbf{P}_{3}=\mathbf{u}_{3}\mathbf{u}_{3}^{\mathsf{T}}$. The spectral distributions are set as $\pi_{1}=\mathcal{N}(0.1,1)$, $\pi_{2}=\mathcal{N}(-0.2,1)$, and $\pi_{3}=\mathcal{N}(0.3,1)$. We consider the following three sequential tests: smart Matrix, one that follows the recipe (88) with $\varepsilon=\alpha/20$; naive Matrix, one that is agnostic about the real distribution and picks $\mathbf{A}=(d/\alpha)\mathbf{I}$; and Scalar, all with the parameters $\gamma_{n}=n^{-1/2}$ and $\alpha=0.05$. We record the times at which these tests correctly reject the null, as well as the evolution of $\operatorname{r}^{-1}(\mathbf{Y}_{n}^{\mathsf{SN}})={\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}$. The findings are displayed in Figure 1: smart Matrix indeed outpowers the other two, whereas naive Matrix and Scalar are similar in power, explained by the quick convergence of $\operatorname{r}(\mathbf{Y}_{n}^{\mathsf{SN}})$ to 1.
Figure 1: Empirical findings with commuting i.i.d. random matrices. Left. Comparison of the cumulative distribution functions of rejection times of the sequential tests corresponding to the matrix- and scalar-valued test processes under the same alternative distribution. Right. Evolution of the inverse of the effective rank of the matrix-valued test process under the alternative.
@@
which indeed satisfies $(\mathrm{tr}\mathbf{\Sigma})^{2}+2\mathrm{tr}(\mathbf{\Sigma}^{2})=\frac{61}{144}<1$. The three sequential tests as before are compared in Figure 2: the smart Matrix that follows (90) with $\varepsilon=\alpha/20$; the naive Matrix with $\mathbf{A}=(d/\alpha)\mathbf{I}$; Scalar, all with $\gamma_{n}=n^{-1/2}$ and $\alpha=0.05$. Besides the unsurprising fact that the smart Matrix is again the most powerful, rejecting the null earlier than the other two, we also observe the phenomenon that, while in the $n=1$ case, the matrix-valued test process with “naive” $\mathbf{A}=(d/\alpha)\mathbf{I}$ is always less powerful than the scalar-valued test process, this is reversed under non-commutativity with large $n$ as naive Matrix rejects slightly earlier than Scalar in our simulation. As also shown in Figure 2, the combination of two auxiliary findings contributes to this phenomenon: $\operatorname{r}(\mathbf{Y}_{n}^{\mathsf{SN}})={\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}\to 1$ even when observations do not commute; and, while $L_{n}^{\textsf{SN}}=\mathrm{tr}(\mathbf{Y}_{n}^{\textsf{SN}})$ in all our previous examples, now $\mathrm{tr}(\mathbf{Y}_{n}^{\textsf{SN}})$ grows slightly faster than $L_{n}^{\textsf{SN}}$. It takes further theoretical investigations to understand why this happens (either in this particular example or in general), but to some extent it hints nonetheless at the possibility that in practice where the effect size is small (requiring large $n$) and matrix observations are non-commuting, there is an inherent advantage using the matrix-valued test process we propose.
Figure 2: Empirical findings with non-commuting i.i.d. random matrices for covariance testing. Left. Comparison of the cumulative distribution functions of rejection times of the sequential tests corresponding to the matrix- and scalar-valued test processes under the same alternative distribution. Middle. Evolution of the inverse of the effective rank of the matrix-valued test process under the alternative. Right. Comparison between the trace of the matrix-valued test process, and the scalar-valued test process under the same alternative distribution.
@@ 7 Conclusion
In this paper, we study positive semidefinite matrix supermartingales and backward submartingales. In particular, we present several maximal inequalities that generalize existing results in the well-known scalar case. We demonstrate the statistical applications of these results by noting that for data and hypotheses that inherently take matrix values, our “matrix-native” approach is beneficial and can outperform previous methods that often require “scalarization”.
In particular, sequential testing with matrix-valued test supermartingales and e-processes allows a freely chosen rejection threshold matrix $\mathbf{A}$ subject to the constraint $\mathrm{tr}(\mathbf{A}^{-1})=\alpha$, one that can be chosen according to the prior knowledge to boost power without sacrificing the type 1 error guarantee. On the other hand, taking the threshold matrix $\mathbf{A}$ to be a multiple of the identity matrix, a degenerate choice that it appears, can still lead to a more powerful test in the large sample regime compared to scalarization. This is an empirical finding of our paper that calls for further understanding.
Our study shall lead to more powerful and flexible sequential hypothesis testing procedures for matrix inference problems including covariance and graph testing.
@@ Acknowledgement
The authors thank Joel A. Tropp, Lester Mackey, Tudor Manole, and Reihaneh Malekian for suggestions and discussions, and acknowledge support from NSF grant DMS-2310718.
Appendix A On the Randomized Matrix Markov Inequality
Instead of randomizing the bound with the isotropic diagonal matrix $U\mathbf{I}$ where $U$ is a uniform or super-uniform random scalar, let us seek an inherently matrix-born definition of uniformity and super-uniformity. Scalar super-uniformity has two viable extensions to matrices, one in terms of the trace and the other of the spectral norm.
Definition A.1.
Straightforwardly the identity matrix $\mathbf{I}$ is spectrally (hence trace) super-uniform. If $U\sim\operatorname{Unif}_{(0,1)}$, $U\mathbf{I}$ is $\mathcal{S}_{d}^{[0,1]}$-spectrally uniform, and $\left\{\mathbf{v}\mathbf{v}^{\mathsf{T}}:\|\mathbf{v}\|\leqslant 1\right\}$-trace uniform. Having defined trace super-uniformity, we can now state a matrix extension of the uniformly randomized Markov’s inequality (72).
Theorem A.2 (Generalized Uniformly Randomized Matrix Markov Inequality).
$\mathbb{P}(\mathbf{X}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{X})\mathbf{A}^{-1}).$ (92)
Proof.
By our assumptions,
@@ Appendix A On the Randomized Matrix Markov Inequality
concluding the proof. ∎
When taking $\mathbf{U}$ to be the deterministic $\mathbf{I}$, (92) recovers (MMI). It is not hard to see that (92) is strictly more general than the randomized scalar Markov’s inequality (72) and the scalar-randomized matrix Markov’s inequality (73). We also see from the monotonicity of the trace and the trace-spectral norm inequality that scalarized trace and spectral bounds are easily obtainable via
The following is an example where the inequality of (92) holds with equality. We take $\mathbf{U}=U\mathbf{I}$ where $U\sim\operatorname{Unif}_{(0,1)}$, and $\mathbf{A}\in\mathcal{S}_{d}^{++}$. Let $\mathbf{x}$ be a random vector in the ellipsoid $\{\mathbf{x}:\mathbf{x}^{\mathsf{T}}\mathbf{A}^{-1}\mathbf{x}\leqslant 1\}\subseteq\mathbb{R}^{d}$. Then,
where the second equality above is (92) applied on the random matrix $\mathbf{x}\mathbf{x}^{\mathsf{T}}$. It is not hard to see here the condition for equality is met. Intuitively, this hints that (92) is tighter when the random matrix $\mathbf{X}$ is approximately of lower rank, a point that we shall return to later in this paper with statistical examples.
Appendix B Omitted Proofs
B.1 Dominated Integrability Lemma